OpenStudy (anonymous):

7r-5r+2=5r-r step by step plz

jimthompson5910 (jim_thompson5910):

7r-5r+2=5r-r 2r+2=5r-r 2r+2=4r 2r=4r-2 2r-4r=-2 -2r=-2 r=-2/(-2) r=1

OpenStudy (mathteacher1729):

For problems like these... go to http://www.wolframalpha.com type solve 7r-5r+2=5r-r click "show steps" :)